I am having a word problem issue and was hoping you could assist me.The Problem is:
Someone invested $7500 into two mutual funds. One fund earned 4% profit while the other took a loss of 13% Despite the loss, this person made a profit of $45. how much did he invest in both funds.
Can you help me solve this problem?

Answer by rothauserc(4718)

let x be the amount invested in the 4% profit fund and y be the amount invested in the 13% loss fund, then
x + y = 7500
.04x -.13y = 45
solve first equation for x and substitute for x in second equation
.04(7500-y) -.13y = 45
300 -.04y -.13y = 45
-.17y = -255
y = 1500
x = 6000
$6000 was invested at 4% and $1500 was invested at a loss of 13%
******************************************
now check the answer by substituting in second equation
.04(6000) -.13(1500) = 45
240 - 195 = 45
45 = 45
our answer checks